OpenStudy (anonymous):

the logarithms of 2 to the base 16 =y solve for y

OpenStudy (amistre64):

logB(arg) = exp log16(2) = y perhaps? means 16^y = 2

OpenStudy (anonymous):

ya i got 16^y=2 but how to you get y

OpenStudy (amistre64):

16^1 = 16 16^0 = 1 so its between 0 and 1 :)

OpenStudy (amistre64):

logs and exponents are inverse; you can try to see if there is a 2^n that equals 16 to help out

OpenStudy (amistre64):

then flip it and see if it works :)

OpenStudy (amistre64):

16^(1/2) = sqrt(16) = 4 so its less than 1/2

OpenStudy (amistre64):

sqrt(4) = 2 tho sqrt(sqrt(16)) = 16^(1/2)^(1/2) which equals 16^(1/4)

OpenStudy (anonymous):

it is..

OpenStudy (anonymous):

16^y=2 (4^2)^y=2^1 2y=1 y=0.5

OpenStudy (amistre64):

\[B^{n^m} = B^{n*m}\]

OpenStudy (amistre64):

.5 is close, but not quite B^(1/2) = sqrt(B) 16^(1/2) ?=? 2 sqrt(16) ?=? 2 4 ?=? 2 nope

OpenStudy (anonymous):

no that is the answer lol.. its in my book

OpenStudy (anonymous):

nvm i wrote it wrong

OpenStudy (anonymous):

the answer is..

OpenStudy (amistre64):

:) playing games are we lol

OpenStudy (anonymous):

16^y=2 (2^4)^y=2^1 4y=1 y=0.25

OpenStudy (amistre64):

thats correct

OpenStudy (anonymous):

yusss!

OpenStudy (anonymous):

how about log base x 20=1

OpenStudy (amistre64):

logX(20) = 1 means X^1 = 20 x = 20 then right?

OpenStudy (anonymous):

k good.. checking my work:P thanks!

OpenStudy (amistre64):

logB(B) = 1 is a standard rule of logs

OpenStudy (anonymous):

how about log base x x^5=5

OpenStudy (amistre64):

logx(x^5) = 5 ; logB(M^n) = n logB(M) 5 logx(x) = 5 ; logB(B) = 1, so x is all values that are positive

OpenStudy (anonymous):

huh?

OpenStudy (amistre64):

plug in any value for x and you get the same results; there is no 1 solution, it is infinite

OpenStudy (anonymous):

it will always equal 5?

OpenStudy (amistre64):

barring of course for the fact that log(n<=0) is a nono

OpenStudy (amistre64):

any value of x will work so long as its greater than 0

OpenStudy (amistre64):

log100(100^5) = 5 log2(2^5) = 5 log1/23(1/23^5) = 5

OpenStudy (anonymous):

how about.. log base 32 y =2/5

OpenStudy (amistre64):

log32(y) = 2/5 y = 32^(2/5) then

OpenStudy (amistre64):

logB(y) = x means y = B^x

OpenStudy (anonymous):

does y = 4?

OpenStudy (amistre64):

dunno, id have to calculator it :)

OpenStudy (amistre64):

yes y=4

OpenStudy (anonymous):

i wrote 4 so maybe ill keep it like that.. how about.. log y 81=4/3 :)

OpenStudy (amistre64):

post some new qs so others can get a chance :) and good luck, youre doing great!!

OpenStudy (anonymous):

ahh but you're the best!